OpenStudy (beaubrun):

College Algebra Questions 2? 1. Determine, algebraically, whether the function is even, odd, or neither even nor odd? f(x) = x^3 + x^2 - 2 2. Use the given conditions to write an equation for the line in slope-intercept form? Passing through (-3, -6) and (-5, -7) 3. Find and simplify the difference quotient f(x + h) - f(x)/h, h ≠ 0 for the given function? f(x) = x^2 + 4x - 7 4. Find the average rate of change of the function from x1 to x2? f(x) = 2x from x1 = 2 to x2 = 8

OpenStudy (john_es):

A function is even if, \[f(-x)=f(x)\] A function is odd if, \[f(-x)=-f(x)\]

OpenStudy (john_es):

For the second, if you have two points, you can find the slope-intercept, \[y-y_1=\frac{y_2-y_1}{x_2-x_1}(x-x_1)\]

OpenStudy (john_es):

And for the third, the average rate of a function between x_1 and x_2, \[\Delta f=\frac{f(x_2)-f(x_1)}{x_2-x_1}\]

OpenStudy (john_es):

Sorry that was the fourth.